Fermi National Laboratory


Accelerator Update for January 19, 2005 - January 21, 2005

Wednesday January 19
The day shift began with Operations monitoring store 3926 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson experiments MCenter and MTest taking beam.

Operations reported at 10:21 AM that the Linac I- Source was suffering from regulation problems. Experts investigated and found a problem with the water-resistor water-skid.

Operations began stashing antiprotons into the Recycler at 12:30 PM. There were nine transfers in all.

A Linac expert turned off the I- Source at 1:19 PM to make repairs. This shut off all the downstream beam. The expert finished his work and operators had beam restored to MiniBooNE by 1:43 PM.

The Linac expert reported at 1:58 PM that the water-resistor was dirty. It would be cleaned by back-flushing during the Linac access on Thursday.

Pbar resumed stacking at 2:12 PM.

An expert reported at 2:23 PM that he had replaced the power supply for a Booster quadrupole (QDB3).

Operations established beam to the Switchyard (SY) at 2:33 PM.

MCenter reported at 5:24 PM that they had to turn off due to the beam intensity fluctuations.

At 5:43 PM, store 3926 aborted. Operators believed that a controls hardware failure caused the abort.

Operators prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access at 9:08 PM. The experimenters were out of access by 11:36 PM.

Operations put the Tevatron (TeV) into a dry squeeze at 11:58 PM.

Thursday January 20
The midnight shift began with the TeV conducting studies, with Operations monitoring the stash, and with MiniBooNE, MCenter, and MTest taking beam.

Operations put the TeV through another dry squeeze at 2:02 AM.

Operations put the TeV into shot setup at 2:33 AM. (Any shot setup, whether it's for the TeV or Recycler, halts SY beam.)

Operations established store 3929 at 5:10 AM with an initial luminosity of 90.74E30. Pbar resumed stacking.

Operations turned off Linac at 6:31 AM to make ready for the eventual accesses into Linac, Booster, MiniBooNE, and the Main Injector.

Operators began preparing MI for access at 6:42 AM.

Operators turned off Booster at 8:45 AM.

MiniBooNE came out of access at 11:07 AM.

Linac came out of access at 11:44 AM.

At 1:40 PM, an MIRF expert reported that the RF LCW leak would take another hour to a hour and a half to repair.

Booster came out of access and an expert began tuning beam to the dump at 2:15 PM.

At 2:42 PM, Operations reported that they lost all safety system links to the interlock systems which in turn caused the loss of store 3929 and the Pbar stack. Safety system experts investigated and discovered that an inadvertent short while installing new equipment had caused the failure.

Operators prepared the TeV sector F for access at 3:13 PM to repair a kicker. Operators also prepared CDF and D0 for access.

Operations reported that the Neutron Therapy Facility (NTF) began taking beam at 3:50 PM. They took beam for about an hour and a half and then shut off. With this beam they resumed patient treatment.

Cryo system technicians began work at 4:08 PM on a TeV sector E wet engine.

The TeV F sector work finished at 4:45 PM.

Operations received permission to send beam to NuMI at 5 PM.

CDF and D0 were out of access by 6 PM.

MI RF techs reported at 7:35 PM that they had completed their work on MIRF11.

Operations established beam to MiniBooNE, MCenter, and MTest at 9:31 PM.

Operations established beam to Pbar at 9:58 PM, but the MI beam wasn't very good. Operators investigated.

Operations put some beam into the TeV for a study period at 10:14 PM.

Friday January 21
The midnight shift began with Pbar stacking, with the TeV conducting studies, and with MiniBooNE, MCenter, and MTest taking beam.

At 1:34 AM, a MI RF station (MIRF18) tripped off for the first of many times throughout the midnight shift.

At 1:54 AM, the TeV aborted due to two bad repeaters. Experts replaced the repeaters.

MTest experimenters reported at 2:52 AM that they were done with the beam for the night.

At 4:46 AM, a quadrupole power supply (I:Q703) tripped off and wouldn't reset. Status indicated that there was a problem with the power supply's VCB. Experts came in and managed to get the supply back on at 5:48 AM without having to go into the VCB.

Operations put the TeV through a dry squeeze at 7:26 AM.

Operations put the TeV into shot setup at 7:50 AM.

The Plan for Friday and the Weekend
The plan for today is to establish a store this morning and then around 10:30 AM or so, start sending beam to NuMI. This weekend the complex will be involved in stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/21/2005   email Fermilab